Eight years from now, clifford will be twice his age was six years ago. What is clifford's present age?
Answer by mananth(16946)


Eight years from now, clifford will be twice his age was six years ago. What is clifford's present age?
Let Clifford's age be x years now
Six years ago (x-6) years
Eight years hence
x+8 years
x+8 = 2(x-6)
x+8 = 2x-12
x=20
x = 20 age now